OpenStudy (anonymous):

Need help! Solve for x: |x| − 8 = −5

OpenStudy (anonymous):

x =3

OpenStudy (anonymous):

x - 8 = -5 x - 8 + 8 = -5 + 8 x = -5 + 8 x = 3

OpenStudy (anonymous):

:)